OpenStudy (anonymous):

The surface area of a cube is 294 square inches. Find the length of a lateral edge. Guys will you please help me? Will medal and fan!

OpenStudy (here_to_help15):

A cube has six faces. Each face has area = 294/6 = 49 Each edge = √49 = 7 inches

OpenStudy (jdoe0001):

hmmm how do you find the area of a square? or a rectangle for that matter?

OpenStudy (here_to_help15):

So since it has 6 sides you would multiply the square inch then etc from there you would \[\sqrt{ }\] it

OpenStudy (anonymous):

each flat area of a cube is L*L there are 6 of these so 6L^2=294 L^2=294/6=49 L=sq.rt.49=7 each edge

OpenStudy (anonymous):

Thank you guys!!! I can only medal one though... But I will fan all of you guys! Thanks my lifesavers!